# Multi-messenger signatures of a deformed magnetar in gamma-ray bursts

### v1

@@ Abstract
We study the evolution of a newly formed magnetized neutron-star (NS) as a power source of gamma-ray bursts (GRBs) in the light of both gravitational wave (GW) and electromagnetic (EM) radiations. The compressible and incompressible fluids are employed in order to model the secular evolution of Maclaurian spheroids. It is shown that the GW and EM light curves evolve as a function of eccentricity and rotational frequency with time. We find that the light curve characteristics crucially depend on NS parameters such as magnitude and structure of magnetic field, ellipticity and the equation of state (EOS) of the fluid. The presence of X-ray flares, whose origins are not yet well understood, can be captured in our model regarding some specific nuclear EOSs. Our model allowing us to explain flares that occur within the wide range of $10$ to $10^{4}$ seconds and the peak luminosity in the order of $10^{46}$ $10^{51}$ $\rm\text{erg}/s$ by using a reasonable set of parameters such as magnetic field strength around $10^{14}-10^{16}$ Gauss, the quadrupole to dipole ratio of magnetic field up to 500. By applying our model to a sample of GRB X-ray flares observed by Swift/XRT, we try to constraint the crucial parameters of a deformed magnetar via MCMC fitting method. Our analysis shows that ongoing and upcoming joint multi-messenger detections can be used to understand the nature of GRB’s central engine and its evolution at the early times of the burst formation.
@@ 1 Introduction
The physics behind GRB’s central engine is still among astrophysics open questions, while many scenarios have been introduced to explain it (e.g., Li et al., 2018). The collapse of a massive compact star is a candidate for long duration GRB’s central engine Lü_2015 and compact binary mergers can be served as a source of short duration GRBs Song & Liu (2023). There are several evidences that the GRB and its afterglow may be powered by a rapidly rotating highly-magnetised NS in which lose of angular momentum is a result of EM and GW emissions or is due to the accretion processes Rowlinson et al. (2013); Mösta et al. (2020); Bucciantini (2012); Bianco et al. (2024). The plateau phase in X-ray afterglow emission is observed in many GRBs (e.g., Liang et al., 2007; Li et al., 2012, 2015; Dainotti et al., 2020, 2022, and references therein) and can be explained by the spin-down of an ultra-magnetized millisecond NS Stratta et al. (2018); Gompertz et al. (2013); Wang et al. (2024). It has been demonstrated that the spin-down luminosity from multipolar EM emission can effectively account for the plateau phase in the majority of GRBs afterglows offering the complex magnetic field structure of magnetars as one of the candidates for GRB’s central engine Wang et al. (2024).
X-ray flares are also frequent events occurring in about 33 percent of GRBs within a time interval ranging from 30 to $10^{5}$ seconds after GRB trigger time. The appearance of early X-ray flares during the light curves has been widely considered previously Ruffini et al. (2018b); Chincarini et al. (2010); Zheng et al. (2022); Saji et al. (2023); Zhang et al. (2006). It has been suggested that the prolonged activity of the central engine in GRB is responsible for flaring features alongside afterglow emissions Burrows et al. (2005); Fan & Wei (2005); Beniamini & Kumar (2016). However explaining the late time flaring activities is challenging in terms of black hole (BH) central engine and are more in favour of the idea of rapidly rotating magnetar central engines for GRBs Rowlinson et al. (2010); Metzger et al. (2008); Zhang & Mészáros (2001); Dai & Lu (1998). It has been investigated that the magneto hydrodynamic instabilities in a rotating neutron star can produce a strong temporary magnetic field of order $10^{17}G$, and the magnetic re-connection effects leads to multiple X-ray flares Kluźniak_1998; Dai et al. (2006). Beyond gamma-ray and X-ray observations, there are various data from optical to radio wavelengths which supports the magnetar central engine model Rowlinson et al. (2013); Xie et al. (2020); Lazarus et al. (2012). These observations offer high values of the magnetic field of the order of $B\sim 10^{12}G-10^{17}G$ with multipolar structures Wang et al. (2024). High values of quadrupole magnetic field component up to $\sim 10^{3}$ times greater than dipole magnetic field have been proposed to explain GRB features Ruffini et al. (2018a); Stratta et al. (2018).
@@
The multi-messenger nature of magnetars through GW observations provides constraints on the deformation parameter, ellipticity, to be of order $\sim 10^{-5}-10^{-2}$asky & Glampedakis (2016). The GW signal generated by magnetars can be explored more precisely by current and upcoming GW detectors Lai & Shapiro (1995); Sur & Haskell (2021); Xie et al. (2022a). These signals are generated primarily due to the extreme magnetic fields and rotational instabilities that cause deviations from axisymmetry. Such observations will provide valuable insights into the physics of dense matter and the strong gravitational fields characteristic of these highly magnetized NSs. The polarization characteristics of the EM counterpart of GW created by coalescence of the binary sources in short GRBs has been considered previously in Shakeri & Allahyari (2018).
@@
In this paper, we consider the evolution of a deforemed magnetar as a power source for GRBs. We use maclaurin spheroid to model a newborn magnetar to explain different properties of both EM and GW observations of GRBs. Maclaurin spheroid as a homogeneous rotating object was first introduced by S.Chandrasekhar (1969) and later on used to model the spin down of pulsars by Finn & Shapiro (1990), where the the evolution of eccentricity leads to spin variation of NS. Despite previous studies Xie et al. (2022a); Lasky & Glampedakis (2016); Suvorov & Kokkotas (2020); Rueda et al. (2022) in which the time evolution of angular velocity is derived directly from rotational energy loss of magnetar, here, we examine angular momentum as a function of the Maclaurin spheroid’s eccentricity, which evolves over time. Our approach is more fundamental since rotational evolution is originating from the deformation of the new NS which can result in both spin down and spin up of the object. It is demonstrated that in the absence of accretion, during the loss of angular momentum, new born NS can also experience spin-up Finn & Shapiro (1990), it is argued that this process eventually leads to the formation of a BH. However, in this paper, we examine the stability regions within the parameter space more accurately where the spin-up is caused by the decreasing moment of inertia due to the softness of the EoS, and BH formation does not occur.
@@
We investigate the impact of key model parameters such as the magnetic field, quadrupole to dipole ratio, mass and EOS on both EM and GW light curves. We assume that both dipole and quadrupole emissions contribute to the EM radiation of the magnetar. We evaluate different features of the X-ray light curves under influence of the properties of the system such as the total mass of the progenitor, the magnetic field of the newborn magnetars for both compressible and incompressible fluids. Moreover, in the case of compressible fluids, we consider the impact of various polytropic (adiabatic) indices on the luminosity features. We see that for some specific values of the polytropic index a peak in the luminosity appears which will be faded out with time, this can provide an explanation for the X-ray flares in GRBs.
@@
The present paper is organized as following. In Sec. 2, we review the main features and properties of the maclaurin spheroid in the context of both incompressible and compressible fluids. Then, we consider the evolutionary equation for the eccentricity due to the variation in angular momentum for both types of mentioned fluids. In Sec. 3, we discuss the evolution of NS’s dipole, quadrupole and total EM luminosity due to evolution of eccentricity caused by angular momentum loss. In Sec. 4, we evaluate the GW radiations of a new born NS and examine the detectability of resulting GW emission in the light of current and future detectors. In Sec. 5, we present total energy loss of the magnetar during spin down due to the evolution of the eccentricity and finally In Sec. 7, we discuss our results and give some concluding remarks. In this paper, we are working in geometrical unit in which $G=c=\hbar=1$.
@@ 2 FRAMEWORK
In this section, we present maclaurin spheroid as a model for the central engine of GRBs. The angular velocity of maclaurin spheroid is a function of eccentricity which measures the deformation of NS in polar plane and is defined as S.Chandrasekhar (1969); Finn & Shapiro (1990) :
@@
where a and c are shown in Fig.1 with c coinciding the rotational axis of neutron star. In order to consider GW emissions, we assume an triaxial configuration of NS with small deformation in equatorial plane so that ellipticity is defined as
@@
with equatorial semi axes a and b shown in Fig.1. For a spheroid with density $\rho$ angular velocity is given by:
@@
$g(e)=\frac{(1-e^{2})^{\frac{1}{2}}}{e^{3}}\arcsin(e)-\frac{3(1-e^{2})}{e^{2}}.$ (4)
The Maclaurin spheroid has angular momentum and mass defined as:
@@
where I is the moment of inertia, here due to the smallness of $\epsilon$ we take $a\approx b$. It was shown that for a equilibrium configuration with constant mass, entropy and chemical potential, the variation of energy, E, is related to the variation of angular momentum, J, according to Ostriker & Gunn (1969); Finn & Shapiro (1990) :
@@
This relation can interpret the transformation of angular momentum into electromagnetic and gravitational energy emissions. During the spin down, eccentricity evolves from a maximum value to a small value, making the star less oblate. Maximum eccentricity for a Maclaurin spheroid, known as bifurcation point of Maclaurin and Jacobian sequence, is approximately 0.813 S.Chandrasekhar (1969), at this point bar-mode instabilities emerge. To avoid these dynamical instabilities, we consider maximum eccentricity value at bifurcation point.
In the following, we apply two different types of fluids i) incompressible and ii) compressible for the maclaurin spheroid to model magnetars as a central engine of GRBs.
Figure 1: A schematic view of an asymmetric rotating NS which rotates around $\hat{z}\rm-{axis}$, the ellipticity $\epsilon$ and eccentricity e are described by Eqs (1) and (2), respectively.
@@ 2.1 Incompressible model
Here, we introduce the incompressible Maclaurin spheroid and it’s evolutionary equations. In this model, the density of the fluid forming the spheroid is constant, therefore semi major axis, $a$, and moment of inertia, $I$, can be written as a function of eccentricity,
@@
where $a_{0}$ and $I_{0}$ are radius and inertial moment of a spherical object, respectively, assuming constant homogeneous density. According to these equations, the time evolution of the eccentricity induces a temporal changes in both $a$ and $I$. Assuming that mass remains unchanged during the evolution, one can find the following equation governing the variation of total angular momentum $J$ by taking the logarithmic derivative of Eqs (3), (5) and (2.1) :
@@
where $g^{{}^{\prime}}\equiv dg/de$. In next subsection, we discuss a more general form of maclaurin spheroid with polytropic density profile.
@@ 2.2 Compressible model
In order to consider compressible maclaurin spheroids, we take polytropic Equation of State (EoS) as
@@
where $P$ is pressure and $\Gamma=1+1/n$ and $n$ is the polytropic index. By solving Poisson’s equation for a uniformly rotating spheroid in hydrostatic equilibrium and applying proper symmetry and boundary conditions one can get Shapiro & Teukolsky (1983) :
@@
Here $\rho_{0}$ is the constant density of a sphere ($e=0$) with radius $a_{0}$ and $A_{3}(e)$ is a function of the eccentricity which is given by:
@@
Note that in the limit of large $\Gamma$, compressible model of the fluid tends to incompressible model. Here, we have generalized form of set of equations (2.1) for compressible model:
@@
Following the same strategy used for incompressible model and taking logarithmic derivative of Eq. (11) one can have:
@@
Where $A_{3}^{{}^{\prime}}\equiv dA_{3}/de$. By substituting Eq. (14) into the logarithmic derivation of Eqs. (5) regarding $\dot{M}=0$, the following equation is obtained:
@@
Knowing the EM and GW emissions, through Eqs.(5) and (7) one can eliminate $J$ and $\dot{J}$ from Eqs. (9) and (15) and find the evolutionary equation for eccentricity for both compressible and incompressible models.
3 EM radiation and Spin down
In this section, we consider EM radiations from a magnetar to determine the evolution of eccentricity and then find the EM light curve. In subsection 3.1, we focus on dipole radiation, followed by calculations for quadrupole radiation in subsection 3.2, and the (dipole + quadrupole) total EM radiation in subsection 3.3.

### v2

# Multimessenger signatures of a deformed magnetar in gamma-ray bursts

@@ Abstract
We study the evolution of a newly formed magnetized neutron-star (NS) as a power source of gamma-ray bursts (GRBs) in the light of both gravitational-wave (GW) and electromagnetic (EM) radiations. The compressible and incompressible fluids are employed in order to model the secular evolution of stable Maclaurian spheroids. It is shown that the GW and EM emissions evolve as a function of eccentricity and rotational frequency with time. We find that the luminosity characteristics crucially depend on NS parameters such as magnitude and structure of magnetic field, ellipticity and the equation of state (EoS) of the fluid. The presence of X-ray flares, whose origins are not yet well understood, can be captured in our model regarding some specific nuclear EoSs. Our model allowing us to explain flares that occur within the wide range of $10$ to $10^{4}$ s and the peak EM luminosity in the order of $10^{46}$ $10^{51}$ $\rm\text{erg}s^{-1}$ by using a reasonable set of parameters, such as magnetic field strength around $10^{14}-10^{16}$ G, the quadrupole-to-dipole ratio of magnetic field up to 500. By applying our model to a sample of GRB X-ray flares observed by the Swift/X-ray Telescope, we try to constraint the crucial parameters of a deformed magnetar via a Marcov Chain Monte Carlo fitting method. Our analysis shows that ongoing and upcoming joint multimessenger detections can be used to understand the nature of a GRB’s central engine and its evolution at the early times of the burst formation.
@@ 1 Introduction
Gamma-ray bursts (GRBs) are broadly classified, based on their observed duration, into short and long bursts (Kouveliotou et al., 1993), each associated with distinct progenitor systems (Piran, 2004; Zhang, 2019). Long GRBs, with durations exceeding 2s, are generally linked to the collapse of massive stars, often Wolf–Rayet stars, at the end of their evolution (MacFadyen & Woosley, 1998; MacFadyen et al., 2001), and the associated supernova signals are confirmed (Galama et al., 1998; Kelly et al., 2008; Kelly & Kirshner, 2012). Short GRBs, typically lasting less than 2s, are thought to originate from the mergers of binary neutron star (NS) or NS–black hole (BH) systems, driven by gravitational wave (GW) inspirals (Berger, 2014; D’Avanzo, 2015; van Putten & Della Valle, 2023).
[…]
The ability of an NS to act as the central engine of a GRB depends on its capacity to produce ultrarelativistic, baryon-poor jets. Newly formed NSs, especially those created in core-collapse events, face challenges due to baryon loading (Eichler & Levinson, 2000). The high initial temperatures of proto-NSs ($T\sim 10^{10}\,\mathrm{K}$) generate intense neutrino fluxes that ablate baryonic material from the NS surface, increasing the jet’s baryonic mass and inhibiting the formation of high Lorentz factor jets (Fujibayashi et al., 2017). In contrast, BH-powered jets are generally cleaner because energy extraction through mechanisms such as the Blandford–Znajek process occurs in the baryon-poor polar regions (Lee et al., 2000). One possible solution involves strong magnetic fields ($B>10^{14}\,\mathrm{G}$) of magnetars, which can confine the jet to narrow, baryon-poor polar regions, thereby reducing baryonic contamination (Bucciantini et al., 2007; Shankar et al., 2021; Wang et al., 2024). Additionally, the observed jet may be produced after the baryons in the surrounding environment have been cleared. Once the jet is formed, it is initially responsible for opening the enclosing shells by creating a cocoon structure. During this process, the initial jet clears the baryonic material along its path, allowing a late, clean, relativistic jet core to emerge (Lazzati & Begelman, 2005; Metzger et al., 2011).
@@
Observationally, there are several pieces of evidence showing that the GRB and its afterglow may be powered by a rapidly rotating, highly magnetized NS in which loss of angular momentum is a result of EM and GW emissions or is due to the accretion processes (Rowlinson et al., 2013; Mösta et al., 2020; Bucciantini, 2012; Bianco et al., 2024; Sarin et al., 2018). To date, GW 170817/GRB 170817A (Abbott et al., 2017) remains the only observed GW event from a binary NS merger, with multiple perspectives on the nature of its outcome. For example, van Putten & Della Valle (2023) argued that a Kerr BH was ultimately formed, inferred from the descending GW chirp that exceeds the limits of a NS, with a 1.7-second delay between the GW signal and EM emission attributed to the collapse of a hypermassive NS into a BH. Gill et al. (2019) suggested that the remnant collapsed into a BH after emitting blue and red kilonova ejecta. The timeline of the collapse aligns with the delay needed for the relativistic jet to emerge. Yu et al. (2018) propose that a long-lived NS could have formed post-merger, which could explain late-time kilonova emission due to energy injection. The plateau phase in X-ray afterglow emission is observed in many GRBs (e.g., Liang et al., 2007; Li et al., 2012, 2015; Dainotti et al., 2020, 2022, and references therein) and can be explained by the spin-down of an ultra-magnetized millisecond NS Stratta et al. (2018); Gompertz et al. (2013); Wang et al. (2024). It has been demonstrated that the spin-down luminosity from multipolar EM emission can effectively account for the plateau phase in the majority of GRB afterglows offering the complex magnetic field structure of magnetars as one of the candidates for the GRB’s central engine (Wang et al., 2024).
[…]
@@
The multimessenger nature of magnetars through GW observations provides constraints on the deformation parameter, ellipticity, to be of order $\sim 10^{-5}-10^{-2}$ Lasky & Glampedakis (2016). The GW signal generated by magnetars can be explored more precisely by current and upcoming GW detectors Lai & Shapiro (1995); Sur & Haskell (2021); Xie et al. (2022a). These signals are generated primarily due to the extreme magnetic fields and rotational instabilities that cause deviations from axisymmetry. Such observations will provide valuable insights into the physics of dense matter and the strong gravitational fields characteristic of these highly magnetized NSs. In the case of short GRBs, it has been shown that X-ray afterglow observations can be applied to constrain the properties of post-merger remnants, providing essential insights into the nuclear EoS and the GW emissions generated by newly formed NSs (Sarin et al., 2020). Moreover, the polarization characteristics of the EM counterpart of the GW created by coalescence of the binary sources in short GRBs have been considered previously in Shakeri & Allahyari (2018).
@@
In this paper, we consider the evolution of a deformed magnetar as a power source for GRBs. We use the Maclaurin spheroid to model a newborn magnetar to explain different properties of both EM and GW observations of GRBs. Maclaurin spheroid as a homogeneous rotating object was first introduced by S.Chandrasekhar (1969) and later on used to model the spin-down of pulsars by Finn & Shapiro (1990), where the the evolution of eccentricity leads to spin variation of NS. Despite previous studies Xie et al. (2022a); Lasky & Glampedakis (2016); Suvorov & Kokkotas (2020); Rueda et al. (2022) in which the time evolution of angular velocity is derived directly from rotational energy loss of magnetar, here, we examine angular momentum as a function of the Maclaurin spheroid’s eccentricity, which evolves over time. Our approach is more fundamental since rotational evolution originates from the deformation of the new NS, which can result in both spin down and spin up of the object. It is demonstrated that in the absence of accretion, during the loss of angular momentum, new born NS can also experience spin-up Finn & Shapiro (1990), it has been argued that this process eventually leads to the formation of a BH. However, in this paper, we examine the stability regions within the parameter space more accurately, where the spin-up is caused by the decreasing moment of inertia due to the softness of the EoS, and BH formation does not occur.
The spin-up mechanism we discuss differs fundamentally from the FRED-shaped light curves observed in many long GRBs, which are generally interpreted as evidence of the central engine spinning down (see, e.g., van Putten, 2012). In our model, the spin-up typically manifests during the early afterglow phase, appearing at the end of the prompt emission and is associated with the early X-ray flare activity. This is in contrast to the FRED profile, which characterizes the early prompt phase and reflects spin-down behavior.
@@
We investigate the impact of key model parameters, such as the magnetic field, quadrupole-to-dipole ratio, mass and EoS, on both EM and GW emissions. We assume that both dipole and quadrupole emissions contribute to the EM radiation of the magnetar. We evaluate different features of the X-ray light curves under the influence of the properties of the system, such as the total mass of the progenitor, the magnetic field of the newborn magnetars, for both compressible and incompressible fluids. Moreover, in the case of compressible fluids, we consider the impact of various polytropic (adiabatic) indices on the luminosity features. We see that for some specific values of the polytropic index, a peak in the luminosity appears, which will fade with time, can provide an explanation for the X-ray flares in GRBs.
@@
The present paper is organized as follows. In Sec. 2, we review the main features and properties of the Maclaurin spheroid in the context of both incompressible and compressible fluids. Then, we consider the evolutionary equation for the eccentricity due to the variation in angular momentum for both types of fluids mentioned. In Sec. 3, we discuss the evolution of the NS’s dipole, quadrupole, and total EM luminosity due to the evolution of eccentricity caused by angular momentum loss. In Sec. 4, we evaluate the GW radiation of a new born NS and examine the detectability of the resulting GW emission in the light of future detectors. In Sec. 5, we present the total energy loss of the magnetar during spin-down due to the evolution of the eccentricity, and finally, in Sec. 7, we discuss our results and give some concluding remarks. In this paper, we are working in geometrical units in which $G=c=\hbar=1$.
@@ 2 FRAMEWORK
In this section, we present the Maclaurin spheroid as a model for the central engine of GRBs. The angular velocity of Maclaurin spheroid is a function of eccentricity, which measures the deformation of NS in the polar plane and is defined as S.Chandrasekhar (1969); Finn & Shapiro (1990) :
@@
where a and c are shown in Fig. 1 with c coinciding the rotational axis of the NS. In order to consider GW emissions, we assume an triaxial configuration of the NS with small deformation in the equatorial plane so that ellipticity is defined as
@@
with equatorial semi-axes a and b shown in Fig. 1. For a spheroid with density $\rho$, the angular velocity is given by
@@
$g(e)=\frac{(1-e^{2})^{\frac{1}{2}}}{e^{3}}(3-2e^{2})\arcsin(e)-\frac{3(1-e^{2})}{e^{2}}.$ (4)
The Maclaurin spheroid has angular momentum and mass defined as
@@
where I is the moment of inertia. Here, due to the smallness of $\epsilon$, we take $a\approx b$. It has been shown that for a equilibrium configuration with constant mass, entropy, and chemical potential, the variation in energy, E, is related to the variation of angular momentum, J, according to Ostriker & Gunn (1969); Finn & Shapiro (1990) :
@@
This relation can interpret the transformation of angular momentum into EM and gravitational energy emissions. During the spin-down, eccentricity evolves from a maximum value to a small value, making the star less oblate. Maximum eccentricity for a Maclaurin spheroid, known as the bifurcation point of a Maclaurin and Jacobian sequence, is approximately 0.813 S.Chandrasekhar (1969), and at this point, bar-mode instabilities emerge. To avoid these dynamical instabilities, we consider the maximum eccentricity value at the bifurcation point. In fact, oscillations of a perturbed NS have different modes among which f-modes and r-modes are the most susceptible to the Chandrasekhar-Friedman-Schutz (CFS) instability, leading to GW emissions Schutz (1980); Chandrasekhar (1970); Chugunov (2019). The instabilities induced by f-modes can be excited for relatively high rotational speed $\Omega\gtrsim 0.8\ \Omega_{k}$ Friedman et al. (1989), where $\Omega_{k}$ is the angular frequency corresponding to the mass-shedding limit, given by
$\Omega_{k}\approx C\sqrt{\left(\frac{M}{M_{\odot}}\right)\left(\frac{10km}{a}\right)^{3}},$ (8)
where $C=7800s^{-1}$. This approximate frequency is obtained for a set of EoSs, marking the point beyond which the centrifugal force would tear apart the NS Haensel & Zdunik (1989); Cook et al. (1994); Krüger et al. (2021); Andersson (2003). On the other hand, r-mode toroidal oscillations, which are triggered by the Coriolis force, are unstable due to GW radiation even for slowly rotating perfect fluid stars Andersson (1998). While r-modes offer a larger instability parameter space, this window will be narrowed down due to the bulk/shear viscosity at high/low temperatures giving rise to a spin threshold of around the sub-kHz regime Glampedakis & Gualtieri (2018); Kraav et al. (2024). The detectability of the r-mode instability is discussed in Kokkotas & Schwenzer (2016); Caride et al. (2019); Mytidis et al. (2015). It was argued that the estimated rates for these events are not high enough to make detection very likely with second-generation GW detectors.
In the following, we consider the Maclaurin spheroid with two different types of fluids, (i) incompressible and (ii) compressible, to model a newborn magnetars as the central engine of GRBs. We focus on the secularly stable spheroids with $0\leq T/|W|\leq 0.13$, where $T$ and $W$ are the rotational energy and gravitational binding energy, respectively. This condition restricts the rotational frequency to be well below the mass-shedding limit; therefore, f-modes can not be excited in this regime. In the case of r-mode instabilities, in spite of their potential, we will not consider them in this paper.
Figure 1: A schematic view of an asymmetric NS that rotates around $\hat{z}\rm-{axis}$, the ellipticity $\epsilon$ and eccentricity e are described by Eqs (1) and (2), respectively.
@@ 2.1 Incompressible Model
Here, we introduce the incompressible Maclaurin spheroid and it’s evolutionary equations. In this model, the density of the fluid forming the spheroid is constant; therefore, the semimajor axis, $a$, and moment of inertia, $I$, can be written as a function of eccentricity,
@@
where $a_{0}$ and $I_{0}$ are the radius and inertial moment of a spherical object, respectively, assuming constant homogeneous density. According to these equations, the time evolution of the eccentricity induces a temporal changes in both $a$ and $I$. Assuming that mass remains unchanged during the evolution, one can find the following equation governing the variation of total angular momentum $J$ by taking the logarithmic derivative of Eqs. (3), (5) and (2.1) :
@@
where $g^{{}^{\prime}}\equiv dg/de$. In next subsection, we discuss a more general form of the Maclaurin spheroid with a polytropic density profile.
@@ 2.2 Compressible model
In order to consider compressible Maclaurin spheroids, we take polytropic EoS as
@@
where $P$ is pressure and $\Gamma=1+1/n$ and $n$ is the polytropic index. By solving Poisson’s equation for a uniformly rotating spheroid in hydrostatic equilibrium and applying proper symmetry and boundary conditions, one can get Shapiro & Teukolsky (1983)
@@
where $\rho_{0}$ is the constant density of a sphere ($e=0$) with radius $a_{0}$ and $A_{3}(e)$ is a function of the eccentricity, which is given by
@@
Note that in the limit of large $\Gamma$, the compressible model of the fluid tends to the incompressible model. Here, we a have generalized form of the set of Eq. (2.1) for the compressible model:
@@
Following the same strategy used for the incompressible model and taking the logarithmic derivative of Eq. (12) one can have:
@@
where $A_{3}^{{}^{\prime}}\equiv dA_{3}/de$. By substituting Eq. (15) into the logarithmic derivation of Eqs (5) regarding $\dot{M}=0$, the following equation is obtained
@@
$\eta=ae/r$ (17)
Knowing the EM and GW emissions, through Eqs.(5) and (7) one can eliminate $J$ and $\dot{J}$ from Eqs. (10) and (17) and find the evolutionary equation for eccentricity for both the compressible and incompressible models.
3 EM Radiation and Spin-down
In this section, we consider EM radiation from a magnetar to determine the evolution of eccentricity and then find the EM light curve. In subsection 3.1, we focus on dipole radiation, followed by calculations for quadrupole radiation in subsection 3.2, and the (dipole + quadrupole) total EM radiation in subsection 3.3.